# Analytic Langlands correspondence from SoV

### v1

@@ 1 Introduction
The proposal made in [EFK1] has two main ingredients helping to gain analytic control. The first is to work in the Hilbert space framework offered by the natural scalar product on spaces of half-densities on the moduli spaces $\mathrm{Bun}_{G}(C)$ of holomorphic $G$-bundles on $C$. Another crucial ingredient is the use of the Hecke-operators, natural families of integral operators on the spaces of half-densities on $\mathrm{Bun}_{G}(C)$ [EFK2]. As eigenfunctions of the Hecke-operators turn out to be simultaneously eigenfunctions of the Hitchin Hamiltonians, one may start by studying the spectral problem for the Hecke operators. This spectral problem is more easily seen to be well-defined, as the Hecke operators are compact self-adjoint operators, while the Hamiltonians of the quantised Hitchin system are unbounded.
@@
To this aim we shall adopt a strategy known as Separation of Variables ($\mathsf{SoV}$) in the literature on quantum integrable models. It amounts to the construction of a unitary integral transformation mapping the product of solutions to an ordinary differential equation to a solution of the eigenvalue equations of our interest. The ordinary differential equations appearing in this context are closely related to the real opers appearing on one side of the analytic Langlands correspondence. The $\mathsf{SoV}$ method is powerful as it allows one to construct eigenfunctions of the Hitchin Hamiltonians from real opers. One may thereby prove that the condition to correspond to a real oper is not only necessary, but also sufficient for being an eigenfunction.
To carry this program out in practise, it will be crucial to understand how the Hecke operators interact with the integral transformations defined by the $\mathsf{SoV}$ method. This will allow us to prove that the image of a real oper under the $\mathsf{SoV}$-transformation is an eigenfunction of the Hecke operators. It turns out that the Hecke eigenvalue property follows from the observation that the result of a Hecke-modification at a point $z\in C\equiv C_{0,n}$ can be represented by a $\mathsf{SoV}$-transformation associated to the Riemann surface $C^{\prime}:=C\setminus\{z\}\simeq C_{0,n+1}$. Proving this fact turns out to be the main technical issue arising in this context.
@@ 2.1 Quantisation of the Hitchin system
We shall define two Hilbert spaces ${\mathcal{H}}_{i}$, $i=0,1$, of homogeneous and translation invariant functions. The elements of ${\mathcal{H}}_{i}$ are represented by functions $\Psi_{i}$ of n1n-1 complex variables $\bm{x}=(x_{1},\dots,x_{n-1})$ satisfying
$\begin{aligned} &\Psi_{i}(x_{1}+\delta,\dots,x_{n-1}+\delta)=\Psi_{i}(x_{1},\dots,x_{n-1}),\\ &\Psi_{i}(\rho x_{1},\dots,\rho x_{n-1})=|\rho|^{2J_{i}}\Psi_{i}(x_{1},\dots,x_{n-1}),\end{aligned},\qquad i=0,1,$ (2.1)
where
@@
It is well-known that such functions can be used to represent invariants in tensor products of $n$ spherical principal series representations of $\mathrm{SL}(2,{\mathbb{C}})$ if $j_{r}\in-\frac{1}{2}+\mathrm{i}{\mathbb{R}}$, $r=1,\dots,n$. Translation invariance allows one to represent the functions $\Psi_{i}$ by the homogenous functions $\Psi_{i}(0,x_{2},\dots,x_{n-1})$. Away from $x_{n-1}=0$ one may represent such functions in the form
$\Psi_{i}(0,x_{2},\dots,x_{n-1})=|x_{n-1}|^{2J_{i}}\psi_{i}(x_{2}/x_{n-1},\dots,x_{n-2}/x_{n-1}).$ (2.3)
Similar representations exist away from $x_{r}=0$, $r=2,\dots,n-1$. It follows that the functions $\psi_{i}$ on the right side of (2.3) represent sections of line bundles ${\mathcal{L}}_{J_{i}}$, $i=0,1$, of densities on the complex projective space ${\mathbb{P}}^{d}$, d=n3d=n-3, in a particular chart.
We may observe that the Hilbert space structure on $L^{2}({\mathbb{C}}^{d})$ allows us to define a natural norm on spaces of regular homogenous functions $\Psi$. With the help of (2.3) we may represent this norm as
$\lVert\Psi\rVert^{2}=\int_{{\mathbb{C}}^{d}}d^{2d}\bm{z}\;|\psi(\bm{z})|^{2},\qquad\bm{z}=(z_{2},\dots,z_{n-2}).$ (2.4)
We furthermore note that the Hilbert spaces ${\mathcal{H}}_{i}$, $i=0,1$, are canonically isomorphic, with unitary isomorphisms $\mathsf{H}_{\infty}:{\mathcal{H}}_{i}\to{\mathcal{H}}_{i+1}$ being defined by sending $\psi_{i}$ to $\psi_{i+1}$, for $i\in{\mathbb{Z}}/2$, but the representation in terms of homogenous functions differs.
@@
$\mathsf{H}_{r}=\sum_{s\neq r}\frac{\eta_{ab}{\mathcal{D}}_{r}^{a}{\mathcal{D}}_{r}^{b}}{z_{r}-z_{s}},\qquad r=1,\dots,n-1,$
where $\eta^{00}=1$, $\eta^{+-}=\eta^{-+}=-\frac{1}{2}$, and $\eta^{ab}=0$ otherwise. One should keep in mind that only n3n-3 out of the n1n-1 operators $\mathsf{H}_{r}$ are linearly independent. We are here considering a slightly more general family of operators compared with [EFK3], having parameters $j_{r}\in-\frac{1}{2}+\mathrm{i}{\mathbb{R}}$ rather than $j_{r}=-\frac{1}{2}$ as considered in [EFK3].
@@ 2.2 Baxter operators
We shall define the Baxter operators $\mathsf{Q}_{y}$ as the composition $\mathsf{Q}_{y}=\mathbf{H}_{\infty}\cdot\mathbf{H}_{y}$, with the Hecke-operators $\mathbf{H}_{y}:{\mathcal{H}}_{i}\to{\mathcal{H}}_{i+1}$, being defined as
@@
$\mathbf{H}_{y}\Psi=\int_{{\mathbb{C}}}d^{2}x\;\mathbb{H}_{y,x}\Psi,\qquad\qquad y_{r}(x_{r},x,y)=-\frac{z_{r}-y}{x_{r}-x}.$
The main properties of operators $\mathsf{Q}_{y}$ have been proven in [EFK3] for $j_{r}=-\frac{1}{2}$, $r=1,\dots,n$. We note that the proofs in [EFK3] can be generalised almost verbatim to the cases with $j_{r}\in-\frac{1}{2}+\mathrm{i}{\mathbb{R}}$ studied here. The Baxter operators are compact, hence bounded, for all $y\in{\mathbb{C}}\setminus\{z_{1},\dots,z_{n-1}\}$. They are furthermore self-adjoint, and mutually commuting,
$\big[\,\mathsf{Q}_{y}\,,\mathsf{Q}_{y^{\prime}}\,\big]=0,\qquad\forall\;y,y^{\prime}\in{\mathbb{C}}\setminus\{z_{1},\dots,z_{n-1}\}.$
The spectral problem for the family of Baxter operators is therefore well-posed, and the abstract spectral theorem ([EFK3, Theorem 3.17]) asserts existence of a discrete set $\mathcal{S}$} such that
@@
where ${\mathcal{H}}_{\chi}$ are the eigenspaces for fixed eigenvalues $\chi$, represented by functions $\chi(y,\bar{y})$ on $C$ such that $\mathsf{Q}_{y}\Psi_{\chi}=\chi(y,\bar{y})\Psi_{\chi}.$ The spaces ${\mathcal{H}}_{\chi}$ are finite-dimensional for all $\chi\in{\mathcal{S}}$.
[EFK3, Proposition 4.3] clarifies the relation between Baxter operators and Hamiltonians, which can be concisely formulated as the operator differential equation
$(\partial_{y}^{2}+\mathsf{T}(y))\mathsf{Q}_{w}=0,\qquad\mathsf{T}(y)=\sum_{r=1}^{n-1}\bigg(\frac{\delta_{r}}{(y-z_{r})^{2}}+\frac{{\mathsf{H}}_{r}}{y-z_{r}}\bigg),$ (2.8)
where $\delta_{r}=j_{r}(j_{r}+1)$. We refer to [EFK3] for the details. This implies that eigenstates of the Baxter operators are eigenstates of ${\mathsf{H}}_{r}$, $r=1,\dots,n-1$.
@@ 2.3 Formulation of the analytic Langlands correspondence
$t(y)=\sum_{r=1}^{n}\bigg(\frac{\delta_{r}}{(y-z_{r})^{2}}+\frac{E_{r}}{y-z_{r}}\bigg),$ (2.9)
having a finite limit $\lim_{y\to\infty}y^{4}t(y)$. An oper on $C_{0,n}$ is a real oper if it has real monodromy.
@@
An oper admits solutions $\chi$ that are single-valued on $C_{0,n}$ if and only if it is real. The solution $\chi$ is uniquely determined by the oper $\partial_{y}^{2}+t(y)$, and vice-versa.
@@ 3 Construction of the SoV transformation
$d\mu_{\rm Skl}(\bm{y})=d^{2(n-3)}\bm{y}\,\prod_{\begin{subarray}{c}r,s=1\\ r<s\end{subarray}}^{n-1}|z_{r}-z_{s}|^{2}\prod_{\begin{subarray}{c}k,l=1\\ k<l\end{subarray}}^{n-3}|y_{k}-y_{l}|^{2}\prod_{r=1}^{n-1}\frac{1}{|\kappa_{r}(\bm{y})|^{2}},$ (3.11)
@@ 3.1 Fourier-transformation
The Fourier-transformation of $D_{\Psi}$ can be defined in the distributional sense, yielding a homogenous distribution $\widetilde{D}_{\Psi}$. It turns out that the homogenous distribution $\widetilde{D}_{\Psi}$ is regular, meaning that there exists a a homogenous function $\widetilde{\Psi}$ such that $\widetilde{D}_{\Psi}=D_{\widetilde{\Psi}}$. Let $\eta_{\rho}:{\mathbb{R}}_{+}\to{\mathbb{R}}$ be a family of smooth functions satisfying $\eta_{\rho}(x)=1$ for $x<\rho$ and $\eta_{\rho}(x)=0$ for $x>2\rho$.
@@
It follows from Lemma (1) that the Fourier-transformation maps the space ${\mathcal{S}}_{J}$ of smooth sections of ${\mathcal{L}}_{J}$ to the space ${\mathcal{S}}_{-J-d-1}$. The definition of the norm given in (2.4) can be applied to $\widetilde{\Psi}$.
@@ 3.2 Separation of Variables
It will be convenient to use “twisted” versions of the Fourier-transformation, sending $\psi$ to
$(\mathsf{F}_{\bm{j},i}\Psi)(\bm{k}):=|\rho(\bm{k})|^{-2(j_{n,i}+1)}\prod_{r=1}^{n-1}|k_{r}|^{2j_{r}+2}\,\widetilde{\Psi}(\bm{k}),\qquad\rho(\bm{k}):=\sum_{r=1}^{n-1}z_{r}k_{r},$ (3.17)
for $i=0,1$, using the notations $j_{n,i}$ introduced in (2.2). The change of variables defined by
$k_{r}(\bm{y},\rho):=\rho\kappa_{r}(\bm{y}),\qquad\kappa_{r}(\bm{y}):=\frac{\prod_{k=1}^{n-3}(z_{r}-y_{k})}{\prod_{s\neq r}(z_{r}-z_{s})},\qquad r=1,\dots,n-1,$ (3.18)
allows us to define
$\Phi(\bm{y},\rho):=(\mathsf{F}_{\bm{j},i}{\Psi})\big(\bm{k}(\bm{y},\rho)\big),$ (3.19)
It follows easily from (3.15) that $\Phi(\bm{y},\lambda\rho)=\Phi(\bm{y},\rho)$, implying that $\Phi(\bm{y},\rho)$ does not depend on $\rho$, defining a function $\phi(\bm{y})$. We then have
@@
The map $\widetilde{{\mathcal{H}}}_{-J-d-1}\to{\mathcal{H}}_{\rm Skl}$ sending $\widetilde{\psi}$ to the function $\phi$ defined above is unitary.
@@
$d^{2(n-3)}\bm{\lambda}=\frac{d^{2(n-3)}\bm{y}}{|\kappa_{n-1}(\bm{y})|^{2(n-2)}}\prod_{\begin{subarray}{c}r,s=1\\ r<s\end{subarray}}^{n-1}|z_{r}-z_{s}|^{2}\prod_{\begin{subarray}{c}k,l=1\\ k<l\end{subarray}}^{n-3}|y_{k}-y_{l}|^{2}.$ (3.21)
It furthermore follows from (3.19) that
@@
with $\{\mathsf{D}^{(\alpha)};\alpha\in{\mathcal{I}}\}$ being a finite set of differential operators in the variables $k_{r}$, and $f_{\alpha}$, $\alpha\in{\mathcal{I}}$ being holomorphic functions on $C$, we may define a differential operator denoted as $\mathsf{D}(y_{k})$ by substituting $z$ by $y_{k}$ “from the left”, and pullback under (3.18), more precisely defined by
@@
The Fourier-transformation $\widetilde{\mathsf{T}}(z):=\mathsf{F}_{\bm{j},i}\cdot\mathsf{T}(z)\cdot\mathsf{F}^{-1}_{\bm{j},i}$ of $\mathsf{T}(z)$ satisfies
@@
By composing (3.25) with the Fourier-transformations $\mathsf{F}_{\bm{j},i}$, it follows from (3.25) that $\mathsf{S}_{\bm{j},i}$ intertwines the eigenvalue equations for the Hamiltonians ${\mathsf{H}}_{r}$ with the system of decoupled equations
$(\partial_{y_{k}}^{2}+t(y_{k}))\Phi(\bm{y})=0,\qquad k=1,\dots,d.$ (3.26)

### v2

# Analytic Langlands Correspondence from Separation of Variables

@@ 1 Introduction
The proposal made in [EFK1] has two main ingredients helping to gain analytic control. The first is to work in the Hilbert space framework offered by the natural scalar product on spaces of half-densities on the moduli spaces $\mathrm{Bun}_{G}(C)$ of holomorphic $G$-bundles on $C$. Another crucial ingredient is the use of the Hecke operators, natural families of integral operators on the spaces of half-densities on $\mathrm{Bun}_{G}(C)$ [EFK2]. As eigenfunctions of the Hecke operators turn out to be simultaneously eigenfunctions of the Hitchin Hamiltonians, one may start by studying the spectral problem for the Hecke operators. This spectral problem is more easily seen to be well-defined, as the Hecke operators are compact self-adjoint operators, while the Hamiltonians of the quantised Hitchin system are unbounded.
@@
To this aim we shall adopt a strategy known as Separation of Variables ($\mathsf{SoV}$) in the literature on quantum integrable models. It amounts to the construction of a unitary integral transformation mapping the product of solutions to an ordinary differential equation (ODE) to a solution of the eigenvalue equations of our interest. The ODEs appearing in this context are closely related to the real opers appearing on one side of the analytic Langlands correspondence. The $\mathsf{SoV}$ method is powerful as it allows one to construct eigenfunctions of the Hitchin Hamiltonians from real opers. It shows that having a correspondence to a real oper is not only necessary, as shown in [EFK3, EFK4] for arbitrary $n$, but also sufficient for being an eigenfunction.
To carry this program out in practice, it will be crucial to understand how the Hecke operators interact with the integral transformations defined by the $\mathsf{SoV}$ method. This will allow us to prove that the image of a real oper under the $\mathsf{SoV}$-transformation is an eigenfunction of the Hecke operators. It turns out that the Hecke eigenvalue property follows from the observation that the result of a Hecke modification at a point $z\in C\equiv C_{0,n}$ can be represented by a $\mathsf{SoV}$-transformation associated to the Riemann surface $C^{\prime}:=C\setminus\{z\}\simeq C_{0,n+1}$. Proving this fact turns out to be the main technical issue arising in this context.
@@ 2.1 Quantisation of the Hitchin system
We begin by introducing Hilbert spaces ${\mathcal{H}}_{J}$, of homogeneous and translation invariant functions. The elements of ${\mathcal{H}}_{J}$ are represented by functions $\Psi$ of n1n-1 complex variables $\bm{x}=(x_{1},\dots,x_{n-1})$ satisfying
$\Psi(x_{1}+\delta,\dots,x_{n-1}+\delta)=\Psi(x_{1},\dots,x_{n-1}),$ (2.1)
$\Psi(\rho x_{1},\dots,\rho x_{n-1})=|\rho|^{2J}\Psi(x_{1},\dots,x_{n-1}).$
It is well-known that such functions can be used to represent invariants in tensor products of $n$ spherical principal series representations of $\mathrm{PSL}(2,{\mathbb{C}})$ if $j_{r}\in-\frac{1}{2}+\mathrm{i}{\mathbb{R}}$, $r=1,\dots,n$. Translation invariance allows one to represent the functions $\Psi$ by the homogenous functions $\Psi(0,x_{2},\dots,x_{n-1})$. Away from $x_{n-1}=0$ one may represent such functions in the form
$\Psi(0,x_{2},\dots,x_{n-1})=|x_{n-1}|^{2J}\psi(x_{2}/x_{n-1},\dots,x_{n-2}/x_{n-1}).$ (2.2)
Similar representations exist away from $x_{r}=0$, $r=2,\dots,n-1$. It follows that the functions $\psi$ on the right side of (2.2) represent sections of line bundles ${\mathcal{L}}_{J}$ of densities on the complex projective space ${\mathbb{P}}^{d}$, d=n3d=n-3, in a particular chart.
The Hilbert space structure on $L^{2}({\mathbb{C}}^{d})$ allows us to define a natural norm on spaces of regular homogenous functions $\Psi$. With the help of (2.2) we may represent this norm as
$\lVert\Psi\rVert^{2}=\int_{{\mathbb{C}}^{d}}d^{2d}\bm{\xi}\;|\psi(\bm{\xi})|^{2},\qquad\bm{\xi}=(\xi_{2},\dots,\xi_{n-2}).$ (2.3)
Following [EFK3], we shall use ${\mathcal{H}}_{J}$ to define ${\mathcal{H}}_{i}$ for $i\in{\mathbb{Z}}/2$ as ${\mathcal{H}}_{i}={\mathcal{H}}_{J_{i}}$, with $J_{i}$ chosen as
@@
We may note that the Hilbert spaces ${\mathcal{H}}_{i}$, $i=0,1$, are canonically isomorphic, with unitary isomorphisms $\mathbf{K}:{\mathcal{H}}_{i}\to{\mathcal{H}}_{i+1}$ being represented in terms of the functions $\psi$ representing elements of ${\mathcal{H}}_{i}$ via (2.2) as the identity operator, but different representations in terms of homogenous functions. As explained in [EFK3], one may associate ${\mathcal{H}}_{i}$ with the quantised Hitchin systems associated to vector bundles on $C$ of the form $\mathcal{O}\oplus\mathcal{O}(i)$, for $i=0,1$, respectively.
@@
$\mathsf{H}_{r}=\sum_{s\neq r}\frac{\eta_{ab}{\mathcal{D}}_{r}^{a}{\mathcal{D}}_{s}^{b}}{z_{r}-z_{s}},\qquad r=1,\dots,n-1,$
where $\eta^{00}=1$, $\eta^{+-}=\eta^{-+}=-\frac{1}{2}$, and $\eta^{ab}=0$ otherwise. One should keep in mind that only n3n-3 out of the n1n-1 operators $\mathsf{H}_{r}$ are linearly independent. We are here considering parameters of the form $j_{r}=-\frac{1}{2}+\mathrm{i}s_{r}$ with $s_{r}\in{\mathbb{R}}\setminus\{0\}$, corresponding to the cases of principal series representations of $\mathrm{PSL}(2,{\mathbb{C}})$ discussed in [EFK4], excluding the slightly different case $j_{r}=-\frac{1}{2}$ considered in [EFK3].
@@ 2.2 Baxter operators
We shall define the operators $\mathbf{Q}_{y}$ as the composition $\mathbf{Q}_{y}=\mathbf{K}\cdot\mathbf{H}_{y}$, with the Hecke operators $\mathbf{H}_{y}:{\mathcal{H}}_{i}\to{\mathcal{H}}_{i+1}$, being defined as
@@
$\mathbf{H}_{y}\Psi=\frac{1}{\pi}\int_{{\mathbb{C}}}d^{2}x\;\mathbb{H}_{y,x}\Psi,\qquad\qquad y_{r}(x_{r},x,y)=-\frac{z_{r}-y}{x_{r}-x},\qquad$
We refer to the operators $\mathbf{Q}_{y}$ as Baxter operators [^1] This terminology can be motivated by observing that the equation (2.8) plays the role of Baxter’s T-Q-relations in this context, see also the discussion in [EFK4]. We note, however, that the precise use of this terminology adopted here can differ slightly from other appearances in the literature. in order to distinguish the operators $\mathbf{Q}_{y}:{\mathcal{H}}_{i}\to{\mathcal{H}}_{i}$ from their relatives $\mathbf{H}_{y}$ mapping ${\mathcal{H}}_{i}$ to ${\mathcal{H}}_{i+1}$. The main properties of the operators $\mathbf{Q}_{y}$ have been proven in [EFK4, Section 2]. The Baxter operators are compact, hence bounded, for all $y\in{\mathbb{C}}\setminus\{z_{1},\dots,z_{n-1}\}$. They are furthermore self-adjoint, and mutually commuting,
$\big[\,\mathbf{Q}_{y}\,,\mathbf{Q}_{y^{\prime}}\,\big]=0,\qquad\forall\;y,y^{\prime}\in{\mathbb{C}}\setminus\{z_{1},\dots,z_{n-1}\}.$
The spectral problem for the family of Baxter operators is therefore well-posed, and the abstract spectral theorem ([EFK4, Corollary 2.32]) asserts existence of a discrete set $\mathcal{S}$} such that
@@
where ${\mathcal{H}}_{\chi}$ are the eigenspaces for fixed eigenvalues $\chi$, represented by functions $\chi(y,\bar{y})$ on $C$ such that $\mathbf{Q}_{y}\Psi_{\chi}=\chi(y,\bar{y})\Psi_{\chi}.$ The spaces ${\mathcal{H}}_{\chi}$ are finite-dimensional for all $\chi\in{\mathcal{S}}$.
[EFK4, Proposition 3.10] clarifies the relation between Baxter operators and Hamiltonians, which can be concisely formulated as the operator differential equation
$(\partial_{y}^{2}+\mathsf{T}(y))\mathbf{Q}_{y}=0,\qquad\mathsf{T}(y)=\sum_{r=1}^{n-1}\bigg(\frac{\delta_{r}}{(y-z_{r})^{2}}+\frac{{\mathsf{H}}_{r}}{y-z_{r}}\bigg),$ (2.8)
where $\delta_{r}=j_{r}(j_{r}+1)$. We refer to [EFK4] for the details. This implies that eigenstates of the Baxter operators are simultaneously eigenstates of ${\mathsf{H}}_{r}$, $r=1,\dots,n-1$.
@@ 2.3 Formulation of the analytic Langlands correspondence
$t(y)=\sum_{r=1}^{n}\bigg(\frac{\delta_{r}}{(y-z_{r})^{2}}+\frac{E_{r}}{y-z_{r}}\bigg),\qquad E_{r}\in{\mathbb{C}},\quad r=1,\dots,n,$ (2.9)
having a finite limit $\lim_{y\to\infty}y^{4}t(y)$. An oper on $C_{0,n}$ is a real oper if it has real monodromy, meaning that it has monodromy group conjugate to $\mathrm{SL}(2,{\mathbb{R}})$.
@@
An oper admits solutions $\chi$ that are single-valued on $C_{0,n}$ if and only if it is real [Fi, Theorem 3.5]. The solution $\chi$ is determined uniquely up to scaling by the oper $\partial_{y}^{2}+t(y)$, and vice-versa. Here comes a sketch of the argument from [Te, Fi]. A general solution to (2.10) can locally be represented in the form $\chi(y,\bar{y})=\eta(y)\cdot C\cdot(\eta(y))^{\dagger}$, with $\eta(y)$ being the row vector formed out of two linearly independent solutions to the left equation in (2.10), and $C$ being a two-by-two matrix with complex matrix elements. By a change of basis one may bring $C$ to diagonal form $C=\mathrm{diag}(1,C_{22})$, with $C_{22}\in\{-1,0,1\}$. Single-valued solutions having $C_{22}=1$ or $C_{22}=0$ would define metrics with positive or vanishing constant curvature on $C$, leading to contradictions with the Gauss-Bonnet theorem. Single-valuedness in the remaining case $C_{22}=-1$ requires that the monodromy is in $\mathrm{SU}(1,1)$, and therefore conjugate to $\mathrm{SL}(2,{\mathbb{R}})$.
@@ 3 Construction of the SoV transformation
$d\mu_{\rm Skl}(\bm{y})=d^{2(n-3)}\bm{y}\,\prod_{\begin{subarray}{c}r,s=1\\ r<s\end{subarray}}^{n-1}|z_{r}-z_{s}|^{-2}\prod_{\begin{subarray}{c}k,l=1\\ k<l\end{subarray}}^{n-3}|y_{k}-y_{l}|^{2}\prod_{r=1}^{n-1}\frac{1}{|\kappa_{r}(\bm{y})|^{2}},$ (3.11)
@@ 3.1 Fourier transformation
The Fourier transformation of $D_{\Psi}$ can be defined in the distributional sense, yielding a homogenous distribution $\widetilde{D}_{\Psi}$. It turns out that the homogenous distribution $\widetilde{D}_{\Psi}$ is regular, meaning that there exists a a homogenous function $\widetilde{\Psi}$ such that $\widetilde{D}_{\Psi}=D_{\widetilde{\Psi}}$. Let $\eta_{\rho}:{\mathbb{R}}_{+}\to{\mathbb{R}}$ be a family of smooth functions satisfying $\eta_{\rho}(x)=1$ for $x<\rho$ and $\eta_{\rho}(x)=0$ for $x>2\rho$.
@@
It follows from Lemma (1) that the Fourier transformation maps the space ${\mathcal{S}}_{J}$ of smooth sections of ${\mathcal{L}}_{J}$ to the space ${\mathcal{S}}_{-J-d-1}$. The definition of the norm given in (2.3) can be applied to $\widetilde{\Psi}$.
@@ 3.2 Separation of Variables
A key role will be played by the change of variables $k_{r}=k_{r}(\bm{y},\rho)$ defined by
$\sum_{r=1}^{n-1}\frac{k_{r}(\bm{y},\rho)}{y-z_{r}}=\rho\,\frac{\prod_{k=1}^{n-3}(y-y_{k})}{\prod_{r=1}^{n-1}{(y-z_{r})}}.$ (3.17)
In order to describe the induced maps between spaces of functions, we shall use “twisted” versions $\mathsf{F}_{\bm{j},i}$ of the Fourier transformation, sending $\Psi$ to the functions $\Phi$, defined as
$\Phi(\bm{k})=\bigg|\sum_{r=1}^{n-1}z_{r}k_{r}\bigg|^{-2(j_{n,i}+1)}\prod_{r=1}^{n-1}|k_{r}|^{2j_{r}+2}\,\widetilde{\Psi}(\bm{k}),$ (3.18)
for $i=0,1$, using the notations $j_{n,i}$ introduced in (2.4). We may then note the simple identity
$\Phi(\bm{k}(\bm{y},\rho))=\Phi(\bm{\kappa}(\bm{y})),\qquad\bm{\kappa}(\bm{y}):=\frac{1}{\rho}\bm{k}(\bm{y},\rho),$ (3.19)
following from (3.15) and $\sum_{r=1}^{n-1}z_{r}k_{r}(\bm{y},\rho)=\rho$. The identity (3.19) allows us to define
$\mathsf{SoV}({\Phi})(\bm{y}):=\phi(\bm{y}),\qquad\phi(\bm{y}):=\Phi(\bm{\kappa}(\bm{y})).$ (3.20)
We then have
@@
The map $\mathsf{SoV}:{{\mathcal{H}}}_{-J-d-1}\to{\mathcal{H}}_{\rm Skl}$ defined above is unitary.
@@
$d^{2(n-3)}\bm{\lambda}=\frac{d^{2(n-3)}\bm{y}}{|\kappa_{n-1}(\bm{y})|^{2(n-2)}}\prod_{\begin{subarray}{c}r,s=1\\ r<s\end{subarray}}^{n-1}|z_{r}-z_{s}|^{-2}\prod_{\begin{subarray}{c}k,l=1\\ k<l\end{subarray}}^{n-3}|y_{k}-y_{l}|^{2}.$ (3.22)
It furthermore follows from (3.16) and (3.20) that
@@
with $\{\mathsf{D}^{(\alpha)};\alpha\in{\mathcal{I}}\}$ being a finite set of differential operators in the variables $k_{r}$, and $f_{\alpha}$, $\alpha\in{\mathcal{I}}$ being holomorphic functions on $C=\mathbb{P}^{1}\setminus\{z_{1},\dots,z_{n}\}$, we may define a differential operator $\mathsf{D}(y_{k})$ by substituting $z$ by $y_{k}$ “from the left”, and pullback under (3.17), more precisely
@@
The Fourier transformation $\widetilde{\mathsf{T}}(z):=\mathsf{F}_{\bm{j},i}\cdot\mathsf{T}(z)\cdot\mathsf{F}^{-1}_{\bm{j},i}$ of $\mathsf{T}(z)$ satisfies
@@
By composing (3.26) with the Fourier transformations $\mathsf{F}_{\bm{j},i}$, it follows from (3.26) that $\mathsf{S}_{\bm{j},i}$ intertwines the eigenvalue equations for the Hamiltonians ${\mathsf{H}}_{r}$, or equivalently the equations $\mathsf{T}(z)\Psi=t(z)\Psi$, with the decoupled system of equations
$\partial_{y_{k}}^{2}\Phi(\bm{y})=t(y_{k})\Phi(\bm{y}),\qquad k=1,\dots,d.$ (3.27)